How many mols of Pb are in 1.48 x 1025Pb atoms?
1
Expert's answer
2020-10-07T07:33:22-0400

One mole of lead contains 6.023×1023 atoms of lead

6.023×1023atoms = 1mol

1atoms = (1/6.023×1023)mol

-->1.48×1025atoms =

1.48×1025×(1/6.023×1023)mol

         =24.57mol.